OpenStudy (greatbritain00):

How could you solve sin^2x=2sinx??

OpenStudy (greatbritain00):

PLS HELP

OpenStudy (greatbritain00):

Will give medals... Really need help though

OpenStudy (damir):

ALL POSIBLE WAYS

OpenStudy (greatbritain00):

What do you mean by that though?

OpenStudy (damir):

I MEAN ANY WAY YOU DO IT IT WILL BE CORRECT

OpenStudy (damir):

NEED MORE EXPLANATION?

OpenStudy (greatbritain00):

Yes, please. I just couldn't figure out how to start it

OpenStudy (damir):

SO IF WE PUT ANY NUMBER LIKE 1,2,3,4.5 ETC THE PROBLEM WILL WORK

OpenStudy (damir):

IN THE X SPOT

OpenStudy (greatbritain00):

I can see that you solved it, but what were the steps to figure out that the answer is all reals?

OpenStudy (damir):

WELL GIVE ME A MIN TO WRITE

OpenStudy (damir):

SO WHEN YOU HAVE sin^2x=2sinx?? YOU MUST FIGURE WHAT THE X MEANS SO 

OpenStudy (damir):

GOT IT? @greatbritain00

myininaya (myininaya):

they aren't the same thing you can get everything on one side and factor

myininaya (myininaya):

\[\sin^2(x)=2\sin(x) \\ \sin^2(x)-2\sin(x)=0\] factor the left hand side

myininaya (myininaya):

notice both terms have a common factor of sin(x)

OpenStudy (damir):

I KNOW THAT

OpenStudy (greatbritain00):

Oh! I see.

OpenStudy (greatbritain00):

So sinx=0 and sinx=2 Then I just solve for x? Right?

myininaya (myininaya):

\[\sin^2(x)=2\sin(x) \\ \sin^2(x)-2 \sin(x)=0 \\ \sin(x) [\sin(x)-2]=0 \\ \sin(x)=0 \text{ or } \sin(x)=2\] yep good

OpenStudy (greatbritain00):

Thank you so much!

myininaya (myininaya):

remember that sin(x) only can output values between -1 and 1 and 2 is outside the output range so you only need to focus on sin(x)=0

OpenStudy (greatbritain00):

Sounds good! So x=k (pi)

OpenStudy (greatbritain00):

and KEZ

myininaya (myininaya):

awesome

myininaya (myininaya):

lol somehow i knew you meant k is an element of the integers when you wrote KEZ

OpenStudy (greatbritain00):

Great minds think alike